OpenStudy (anonymous):

Solve the equation. Check both solutions and only write the real solution(s). square root 2x+3=x

OpenStudy (anonymous):

x=-3

OpenStudy (anonymous):

sqrt(2x+3)= x square both sides 2x+3= x^2 set it in standard form -x^2 + 2x+3= 0 use the quadratic formula where ax^2 + bx + c = 0 use quadratic formula to solve for x http://en.wikipedia.org/wiki/Quadratic_equation